# On the number of linear uniform hypergraphs with girth constraint

### v1

@@ 5.1 A special injective function
because there are at most $i^{|{U}|}$ ways that all $r$-sets in U{U} appear in $\mathbb{H}(i)$, and each $r$-set is added to $\mathbb{H}(i)$ at some step $j\leqslant i$ with probability at most 1/|Q(i)|1/|{Q}(i)|. Since $p_{i},\xi_{i}=\Theta(1)$ when $t\in[0,t_{{M}}]$ as shown in (3.5) and (3.9), we have $|{Q}(i)|=\Theta(n^{r})$ as the equation shown in (3.19), and it follows that
@@
where $j\in[s]$. By double-counting arguments, for any $s$-tuples of graphs $(\widetilde{H}_{1},\ldots,\widetilde{H}_{s})$ with $\widetilde{H}_{j}$ satisfying $\hat{H}\subseteq\widetilde{H}_{j}\subseteq H$, we have at most $\prod_{j\in[s]}n^{v_{H}-v_{\widetilde{H}_{j}}}$ ways to generate $s$-tuples $(\psi_{1},\cdots,\psi_{s})\in(\Psi_{\theta,\hat{H},H}(i))^{s}$.
@@
$=O\biggl(\Bigl(\max_{\hat{H}\subseteq\widetilde{H}\subseteq H}n^{[v_{H}-(r-1)e_{H}]+[(r-1)e_{\widetilde{H}}-v_{\widetilde{H}}]}\bigl(nt\bigr)^{e_{H}-e_{\widetilde{H}}}\Bigr)^{s}\biggr).$
For any $\epsilon>0$, choosing $s=\bigl\lceil(2+2r\ell+\mu)/\epsilon\bigr\rceil$, by Markov’s inequality of conditional probability, we have
@@
$=O\bigl(n^{-\epsilon s}\bigr)=O\bigl(n^{-(2+2r\ell+\mu)}\bigr).$
Finally, by taking a union bound for all possible $(i,\hat{H},H,\theta)$, which has at most $o(n^{2+2r\ell})$ ways because there are ${M}=o(n^{2})$ ways to choose $i$, and $O(n^{2r\ell})$ ways to choose $(\hat{H},H,\theta)$ as $\max\{v_{H},e_{H}\}\leqslant 2r\ell$, we have $\mathbb{P}[\bar{\mathcal{C}}(i)\,|\mathcal{G}(i)]=o(n^{-\mu})$ to complete the proof of Claim 5.1. ∎
@@ 5.2 View the parameters in a different way
By Claim 5.1, for $1\leqslant i<M$, assuming the event $\mathcal{G}(i)\cap\mathcal{C}(i)$ holds, we introduce some notation to recharacterize the overcounting and boundedness parameters in Claim 4.1.3 and Claim 4.1.6.
@@
In particular, $\boldsymbol{f}_{m}=\boldsymbol{f}=\emptyset$ when $m=0$. Let $\Gamma_{{U},\boldsymbol{f}_{m},L,k}(i)$ be
@@
where $\alpha^{\prime}\in(\alpha_{0},\alpha)$.
@@
For any $\boldsymbol{f}_{m}\in{K}_{m}(i)$ with $2\leqslant m\leqslant r-1$ and $\boldsymbol{g}\in{Q}(i)$, if $|\boldsymbol{f}_{m}\cap\boldsymbol{g}|\leqslant 1$, by the set $\Pi_{\boldsymbol{f}_{m},\boldsymbol{g}}(i)$ defined in Section 4.1.4 and $\Gamma_{{U},\boldsymbol{f}_{m},L,k}(i)$ in (5.6), we have
@@
For any $L\in\mathcal{L}$, substituting ${U}=\{\boldsymbol{g}\}$ and $k=e_{L}-2$ in Lemma 5.4, we have
@@
For any $\boldsymbol{f}$, $\boldsymbol{g}\in{Q}(i)$, $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$, if $\boldsymbol{f}\neq\boldsymbol{g}$, by the set $\Pi_{\boldsymbol{f},L,k,\boldsymbol{g}}(i)$ defined in Subsection 4.1.4 and $\Gamma_{{U},\boldsymbol{f}_{m},L,k}(i)$ in (5.6), we have
@@
With the help of $\Gamma_{{U},\boldsymbol{f}_{m},L,k}(i)$ in (5.6), we further define some sets to collect overlapping copies of the pair $L$ and $K$ for any $(L,K)\in\mathcal{L}^{2}$, which are extended from certain $r$-sets in $Q(i)$.
@@
By Remark 5.6 and Lemma 5.7, note that $e_{L},e_{K}\leqslant\ell$ and $|\mathcal{L}|=O(1)$, we complete the proof of the equations (4.11)-(4.13) in Claim 4.1.3. ∎
@@
By Remark 5.6 and Lemma 5.7, note that $e_{K}\leqslant\ell$ and $|\mathcal{L}|=O(1)$, we complete the proof of the equation (4.23) in Claim 4.1.6. ∎
@@
Consider any $\widetilde{H}$ with $\hat{H}\subseteq\widetilde{H}\subseteq H$. Note that for $t\in[0,t_{{M}}]$ with $t_{{M}}$ in (3.26),
$n^{\epsilon}(nt)^{e_{H}-e_{\widetilde{H}}}\leqslant n^{\epsilon}(nt_{M})^{\ell-2}=o(n^{\alpha^{\prime}})$
in Lemma 5.1 because $\alpha^{\prime}\in(\alpha_{0},\alpha)$ with $\alpha_{0}=\frac{\ell-2}{\ell-1}$ and $\alpha\in(\alpha_{0},1)$ in (3.27). By Lemma 5.1, it suffices to show
@@
Now suppose that $V(\widetilde{H})\subsetneq V(H)$. Since $L$ is a linear cycle of length at most $\ell$ and $v_{L}=(r-1)e_{L}$ by Remark 3.2, $H\subset L$ with $|H|=k$ for $0\leqslant k\leqslant e_{L}-2$, we have the number of edges of $L$ induced on $V(\widetilde{H})\cup\boldsymbol{f}^{\prime}$ is at most $\frac{v_{\widetilde{H}}+\mathbbm{1}_{\{m\geqslant 1\}}(r-m-1)}{r-1}$. Note that $\H{U}^{\prime}\cup\{\boldsymbol{f}^{\prime}\}\subseteq L\backslash$ H and we further delete one from $e_{\widetilde{H}}$ when $|{U}|+\mathbbm{1}_{\{m\geqslant 1\}}<e_{L}-k$. Finally, we have
$e_{\widetilde{H}}\leqslant\frac{v_{\widetilde{H}}+\mathbbm{1}_{\{m\geqslant 1\}}(r-m-1)}{r-1}-|{U}|-\mathbbm{1}_{\{m\geqslant 1\}}-\mathbbm{1}_{\{|{U}|+\mathbbm{1}_{\{m\geqslant 1\}}<e_{L}-k\}}.$
Now $v_{H}-(r-1)e_{H}]-[v_{\widetilde{H}}-(r-1)e_{\widetilde{H}}$] satisfies the inequality shown in (5.12) by replacing $v_{H}=(r-1)e_{L}$ and $e_{H}=k$. ∎
@@
For any $\widetilde{H}$ with $\hat{H}\subseteq\widetilde{H}\subseteq H$, in Lemma 5.1, note that for any $t\in[0,t_{{M}}]$ with $t_{{M}}$ in (3.26)
$n^{\epsilon}(nt)^{e_{H}-e_{\widetilde{H}}}\leqslant n^{\epsilon}(nt_{M})^{k+e_{K}-2}=o\bigl(n^{\alpha^{\prime}}(nt_{{M}})^{k}\bigr),$
where $\alpha^{\prime}\in(\alpha_{0},\alpha)$ with $\alpha_{0}=\frac{\ell-2}{\ell-1}$ and $\alpha\in(\alpha_{0},1)$ in (3.27). It suffices to show
@@
Note that $V(A)\subseteq V(L)$, $\boldsymbol{f}^{\prime},\boldsymbol{h}^{\prime}\subseteq V(A)$ while $\boldsymbol{f}^{\prime}$, $\boldsymbol{h}^{\prime}\notin\bar{L}$, thus we have $e_{A}\leqslant\frac{v_{A}-\mathbbm{1}_{\{v_{A}\neq v_{L}\}}}{r-1}-2$, which implies
@@
because $L$, $K\in\mathcal{L}$, we have the equation shown in (5.11) holds when $r\geqslant 3$. ∎

### v2

@@ 5.1 A special injective function
because there are at most $i^{|{U}|}$ ways that all $r$-sets in U{U} appear in $\mathbb{H}(i)$, and each $r$-set is added to $\mathbb{H}(i)$ at some step $j\leqslant i$ with probability at most 1/|Q(i)|1/|{Q}(i)|. Since $p_{i}$ in (3.5) and $\xi_{i}$ in (3.9) are $\Theta(1)$ when $t\in[0,t_{{M}}]$ as shown in Remark 3.5, we have $|{Q}(i)|=\Theta(n^{r})$ as the equation shown in (3.19), and it follows that
@@
where $j\in[s]$. By double-counting arguments, for any $s$-tuples of graphs $(\widetilde{H}_{1},\ldots,\widetilde{H}_{s})$ with $\widetilde{H}_{j}$ satisfying $\hat{H}\subseteq\widetilde{H}_{j}\subseteq H$, we have at most $\prod_{j\in[s]}n^{v_{H}-v_{\widetilde{H}_{j}}}$ ways to generate one $(\psi_{1},\cdots,\psi_{s})\in(\Psi_{\theta,\hat{H},H}(i))^{s}$.
@@
$=O\biggl(\Bigl(\max_{\hat{H}\subseteq\widetilde{H}\subseteq H}n^{[v_{H}-(r-1)e_{H}]+[(r-1)e_{\widetilde{H}}-v_{\widetilde{H}}]}\bigl(nt\bigr)^{e_{H}-e_{\widetilde{H}}}\Bigr)^{s}\biggr),$
where the second equality is correct because the ways to choose $(\widetilde{H}_{1},\ldots,\widetilde{H}_{s})$ is finite for the fixed $r$-graph pair $(\hat{H},H)$. For any $\epsilon>0$, choosing $s=\bigl\lceil(2+2r\ell+\mu)/\epsilon\bigr\rceil$, by Markov’s inequality of conditional probability, we have
@@
$\leqslant\mathbb{P}\biggl[\Bigl|\bigl(\Psi_{\theta,\hat{H},H}(i)\bigr)^{s}\Bigr|>n^{\epsilon s}\cdot\Bigl(\max_{\hat{H}\subseteq\widetilde{H}\subseteq H}n^{[v_{H}-(r-1)e_{H}]+[(r-1)e_{\widetilde{H}}-v_{\widetilde{H}}]}\cdot\bigl(nt\bigr)^{e_{H}-e_{\widetilde{H}}}\Bigr)^{s}\,\Big|\mathcal{G}(i)\biggr]$
$=O\bigl(n^{-\epsilon s}\bigr)$
$=O\bigl(n^{-(2+2r\ell+\mu)}\bigr).$
Finally, by taking a union bound for all possible $(i,\hat{H},H,\theta)$ of the above inequality, which has at most $o(n^{2+2r\ell})$ ways because there are $n^{2}t_{M}=o(n^{1+\frac{1}{\ell-1}})$ ways to choose $i$ with $t_{M}$ in (3.26), and $O(n^{2r\ell})$ ways to choose $(\hat{H},H,\theta)$ as $\max\{v_{H},e_{H}\}\leqslant 2r\ell$, we have
$\mathbb{P}[\bar{\mathcal{C}}(i)\,|\mathcal{G}(i)]=o(n^{-\mu})$
to complete the proof of Lemma 5.1. ∎
@@ 5.2 View the parameters in a different way
By Lemma 5.1, for $1\leqslant i<M$, assuming the event $\mathcal{G}(i)\cap\mathcal{C}(i)$ holds, we introduce some notation to recharacterize the overcounting and boundedness parameters in Claim 4.1.3 and Claim 4.1.6.
@@
In particular, $\boldsymbol{f}_{m}=\boldsymbol{f}=\emptyset$ when $m=0$. Let $\Gamma_{{U},\boldsymbol{f}_{m},L,k}(i)$ be defined as
@@
where $\alpha^{\prime}\in(\alpha_{0},\alpha)$ and $\alpha_{0}=\frac{\ell-2}{\ell-1}$.
@@
For any $\boldsymbol{f}_{m}\in{K}_{m}(i)$ with $2\leqslant m\leqslant r-1$ and $\boldsymbol{g}\in{Q}(i)$, if $|\boldsymbol{f}_{m}\cap\boldsymbol{g}|\leqslant 1$, by the set $\Pi_{\boldsymbol{f}_{m},\boldsymbol{g}}(i)$ defined in Section 4.1.4 and $\Gamma_{{U},\boldsymbol{f}_{m},L,k}(i)$ defined in (5.6), we have
@@
For any $L\in\mathcal{L}$, substituting ${U}=\{\boldsymbol{g}\}$, $2\leqslant m\leqslant r-1$ and $k=e_{L}-2$ in Lemma 5.4, we have
@@
For any $\boldsymbol{f}$, $\boldsymbol{g}\in{Q}(i)$, $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$, if $\boldsymbol{f}\neq\boldsymbol{g}$, by the set $\Pi_{\boldsymbol{f},L,k,\boldsymbol{g}}(i)$ defined in Subsection 4.1.4 and $\Gamma_{{U},\boldsymbol{f}_{m},L,k}(i)$ defined in (5.6), we have
@@
With the help of $\Gamma_{{U},\boldsymbol{f}_{m},L,k}(i)$ defined in (5.6), we further define some sets to collect overlapping copies of the pair $L$ and $K$ that are extended from certain $r$-sets in $Q(i)$ for any $(L,K)\in\mathcal{L}^{2}$, where $\mathcal{L}^{2}=\{(L,K)\,|L,K\in\mathcal{L}\}$ is the power set of $\mathcal{L}$.
@@
By Remark 5.6 and Lemma 5.7, we have
$|{R}^{+}_{\boldsymbol{f},L,e_{L}-2,\boldsymbol{f},K}(i)|$ $=O\bigl(n^{\alpha^{\prime}+r-3}(nt_{{M}})^{k}\bigr),$
$|{R}^{+}_{\boldsymbol{f},L,e_{L}-2,\boldsymbol{g},K}(i)|$ $=O\bigl(n^{\alpha^{\prime}+r-3}(nt_{{M}})^{k}\bigr),$
$\mathbbm{1}_{\{k\geqslant 1\}}\cdot|{R}^{++}_{\boldsymbol{f},L,k-1,K}(i)|$ $=O\bigl(n^{\alpha^{\prime}+(r-1)(e_{L}-k)-2}(nt_{{M}})^{k-1}\bigr).$
Note that $e_{L},e_{K}\leqslant\ell$ and $|\mathcal{L}|=O(1)$, we complete the proof of the equations (4.11)-(4.13) in Claim 4.1.3. ∎
@@
By Remark 5.6 and Lemma 5.7,
$|{R}^{+}_{\boldsymbol{f},L,k,\boldsymbol{g},K}(i)|=O\bigl(n^{\alpha^{\prime}+(r-1)(e_{L}-k)-r-1}(nt_{{M}})^{k}\bigr).$
Note that $e_{K}\leqslant\ell$ and $|\mathcal{L}|=O(1)$, we complete the proof of the equation (4.23) in Claim 4.1.6. ∎
@@
Consider any $\widetilde{H}$ with $\hat{H}\subseteq\widetilde{H}\subseteq H$. Note that for $t\in[0,t_{{M}}]$ with $t_{{M}}$ in (3.26), there exists $\alpha^{\prime}\in(\alpha_{0},\alpha)$ with $\alpha_{0}=\frac{\ell-2}{\ell-1}$ such that
$n^{\epsilon}(nt)^{e_{H}-e_{\widetilde{H}}}\leqslant n^{\epsilon}(nt_{M})^{\ell-2}=o(n^{\alpha^{\prime}}).$
By Lemma 5.1, it suffices to show
@@
Now suppose that $V(\widetilde{H})\subsetneq V(H)$. Since $H\subset L$ with $|H|=k$ for $0\leqslant k\leqslant e_{L}-2$, we have the number of edges of $L$ induced on $V(\widetilde{H})\cup\boldsymbol{f}^{\prime}$ is at most $\frac{v_{\widetilde{H}}+\mathbbm{1}_{\{m\geqslant 1\}}(r-m-1)}{r-1}$. In order to obtain the value of $e_{\widetilde{H}}$, note that $\H{U}^{\prime}\cup\{\boldsymbol{f}^{\prime}\}\subseteq L\backslash$ H, then we delete $|{U}|+\mathbbm{1}_{\{m\geqslant 1\}}$ from $\frac{v_{\widetilde{H}}+\mathbbm{1}_{\{m\geqslant 1\}}(r-m-1)}{r-1}$. If $|{U}|+\mathbbm{1}_{\{m\geqslant 1\}}<e_{L}-k$, then $\widetilde{H}$ consists of one linear path or several linear paths because $L$ is a linear cycle of length at most $\ell$ by Remark 3.2. Then we further delete $\mathbbm{1}_{\{|{U}|+\mathbbm{1}_{\{m\geqslant 1\}}<e_{L}-k\}}$ from $e_{\widetilde{H}}$. Finally, we have the value of $e_{\widetilde{H}}$ satisfies
$e_{\widetilde{H}}\leqslant\frac{v_{\widetilde{H}}+\mathbbm{1}_{\{m\geqslant 1\}}(r-m-1)}{r-1}-|{U}|-\mathbbm{1}_{\{m\geqslant 1\}}-\mathbbm{1}_{\{|{U}|+\mathbbm{1}_{\{m\geqslant 1\}}<e_{L}-k\}},$
which implies that
$v_{\widetilde{H}}-(r-1)e_{\widetilde{H}}\geqslant(r-1)\cdot(|{U}|+\mathbbm{1}_{\{|{U}|+\mathbbm{1}_{\{m\geqslant 1\}}<e_{L}-k\}})+\mathbbm{1}_{\{m\geqslant 1\}}\cdot m.$
Since $v_{L}=(r-1)e_{L}$ because $L$ is a linear cycle by Remark 3.2, we have $v_{H}=(r-1)e_{L}$. Now $v_{H}-(r-1)e_{H}]-[v_{\widetilde{H}}-(r-1)e_{\widetilde{H}}$] satisfies the inequality shown in (5.10) by replacing $e_{H}$ by $k$, and $v_{\widetilde{H}}-(r-1)e_{\widetilde{H}}$ by the above inequality. ∎
@@
For any $\widetilde{H}$ with $\hat{H}\subseteq\widetilde{H}\subseteq H$, in Lemma 5.1, note that for any $t\in[0,t_{{M}}]$ with $t_{{M}}$ in (3.26), there exists $\alpha^{\prime}\in(\alpha_{0},\alpha)$ with $\alpha_{0}=\frac{\ell-2}{\ell-1}$ such that
$n^{\epsilon}(nt)^{e_{H}-e_{\widetilde{H}}}\leqslant n^{\epsilon}(nt_{M})^{k+e_{K}-2}=o\bigl(n^{\alpha^{\prime}}(nt_{{M}})^{k}\bigr).$
It suffices to show
@@
Note that $V(A)\subseteq V(L)$, we have $e_{A}\leqslant\frac{v_{A}-\mathbbm{1}_{\{v_{A}\neq v_{L}\}}}{r-1}$ by Remark 3.2. Since $\boldsymbol{f}^{\prime},\boldsymbol{h}^{\prime}\subseteq V(A)$ while $\boldsymbol{f}^{\prime}$, $\boldsymbol{h}^{\prime}\notin\bar{L}$, we further have $e_{A}\leqslant\frac{v_{A}-\mathbbm{1}_{\{v_{A}\neq v_{L}\}}}{r-1}-2$, which implies
@@
because $L$, $K\in\mathcal{L}$, we have the equation shown in (5.11) holds when $r\geqslant 3$ to complete the proof of Lemma 5.7. ∎